OpenStudy (anonymous):

what is 53 over 53 az a fraction in its simplest form

OpenStudy (anonymous):

1

OpenStudy (hoblos):

1

OpenStudy (anonymous):

1

OpenStudy (across):

Are you still missing a screw? :)

OpenStudy (anonymous):

Only a few ;)

OpenStudy (anonymous):

what is that one as a faction

OpenStudy (saifoo.khan):

1/1

OpenStudy (anonymous):

1 = 1/1

OpenStudy (anonymous):

are u sure

OpenStudy (saifoo.khan):

1/(0.5+0.5)

OpenStudy (anonymous):

2/2 3/3 4/4

OpenStudy (saifoo.khan):

wait. 5/5 is the answer

OpenStudy (anonymous):

saifoo what school do u go to i think i might no u

OpenStudy (saifoo.khan):

@zaynelover , i can bet you dont know me! ;) :D

OpenStudy (anonymous):

du go college or secondary atleast tell me that plz

OpenStudy (saifoo.khan):

umm, ever heard about A-levels?

OpenStudy (anonymous):

saifoo.khan is actually my father.

OpenStudy (anonymous):

so u go college

OpenStudy (saifoo.khan):

@fewscrewsmissing is still missing some screws. ;) @zaynelover , yes.. With respect to British Ed Sys.

OpenStudy (anonymous):

cn i put my answer down as 53 over 53

OpenStudy (saifoo.khan):

No, that is NOT simplified.

OpenStudy (anonymous):

how about 1 over 1

OpenStudy (saifoo.khan):

1/1 again equals 1. If u had 1 biscuit and u want to divide it among 1 person. How much that person will get? Obviously 1. So, 1/1 is same thing as 1.

OpenStudy (anonymous):

saifoo.khan, were you born in October 86?

OpenStudy (across):

\[\frac{53}{53}=\frac{106}{106}=\frac{13.25}{13.25}=\frac{\pi}{\pi}=\frac{1}{1}=\frac{1234567890}{1234567890}=\frac{\sum_{n=1}^{\infty}1/n^2}{\sum_{n=1}^{\infty}1/n^2}=1\]It's all the same.

OpenStudy (anonymous):

How's the life of a stalkee treating you these days across?

OpenStudy (across):

I still peek out the window every night before going to sleep to make sure there is nobody hiding behind the bushes.

OpenStudy (saifoo.khan):

Lol, you made me born 10 years before my real DOB/

OpenStudy (anonymous):

I'm glad my legacy lingers ;) The grapevine says you're in a relationship now? Which some dude... Andy or something...? Was quite ticked off about? :-P

OpenStudy (across):

Haha! PM

OpenStudy (anonymous):

have any of uze here got a blackberry

OpenStudy (anonymous):

and is u do do u bbm

OpenStudy (across):

^ o.o ?

OpenStudy (anonymous):

I've got a Blackberry. Want my number? I have no friends and am desperately lonely. I don't even know what my ringtone sounds like.

OpenStudy (saifoo.khan):

Androids > Blackberry.